OpenStudy (anonymous):

Two congruent equilateral triangles have sides of 34 mm and 6x-2 mm. Find the value of x.

OpenStudy (anonymous):

can u draw it down plz

OpenStudy (anonymous):

34 = 6x - 2 36 = 6x x = 36/6 x = 6

OpenStudy (anonymous):

Does that come out to the right answer?

OpenStudy (anonymous):

yes it does . thank you (:

OpenStudy (anonymous):

You're welcome!